To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, April 14-April 18, 1997
Date: April 18, 1997
We conducted a 2-day project review for the Navy High Energy Laser
Office on Monday and Tuesday (April 14-15). The first day focussed
on the status of Instrumentation and Controls (I&C), RF window
testing for the SRF systems, and the Injector Test Stand. On
Tuesday the remaining WBS elements were reviewed by touring FEL
hardware that is currently being tested in the Test Lab and Optics
Trailer or installed in the FEL Building.
On April 15-17, Jefferson Lab hosted a workshop sponsored by the
DOD Laser Systems Safety Working groups. J. Cook presented an
overview of the Navy HEL program and G. Neil presented a summary
of the Jefferson Lab FEL program.
The spring meeting of the SURA Board of Trustees was held in Washington
on April 16-17. F. Dylla preseted a status report on the FEL
program and met with representatives from the Vanderbilt and Duke
FEL Laboratories on matters of potential and continuing collaboration.
The cost and schedule summary data for the month of March was
prepared for the monthly report and for presentation at the April
15 Navy project review.
Injector Test Stand
Having completed bake-out of the photocathode gun, we prepared
the gun and its beamline for operation, including reinstalling
the diagnostic hardware in the beamline that had been removed
for the bake. Then, after reinstalling the sulphur hexafluoride
tank and reactivating the nonevaporable getters in the gun, we
began high-voltage (HV) processing of the new ceramic stack and
coated cathode-support tube. As of now (noon, 18 Apr 97) the
voltage has reached 340 kV after about 8 hours of processing,
a better result than obtained in previous HV processing. The
processing procedure is iterative and takes time because turning
up the voltage entails burning through breakdown barriers as they
are successively reached.
We operated the photocathode drive laser was operated in support
of diagnostics development. It operated normally. We will receive
the new variable-speed electro-optical modulator tomorrow (19
High-power radiofrequency (rf) testing of the cryounit successfully
began. Performance of the subject cavity was reverified up to
its quench limit. Its performance is slightly improved from earlier
tests, probably due to a slight difference in ambient static magnetic
field, and the tests indicate cavity performance remains consistent
with its projected 9 MV/m operation. Up to 18 kW rf was introduced
into the waveguide with no performance degradation of the waveguide.
Today (18 Apr 97) we are switching to the second cavity which
is configured with a polyethylene warm window for 8 kW operation.
Regarding ceramic warm windows, we processed one of the bad ceradyne
ceramics by air- firing it for 30 hours to oxidize fully its bulk
material. Tests of the fired ceramic in the ring- resonator test
assembly indicated greatly reduced dielectric losses, signifying
full recovery of its dielectric performance. This is a very encouraging
Regarding the beam-transport system:
Fabrication of the optical-chicane dipole magnets continued unabated.
Our WBS manager, G. Biallas, is visiting Everson Electric today
(18 Apr 97) to discuss the fine points of the first-article core.
This vendor has completed rough machining and heat treatment.
Magnet Enterprises International completed the first-article
coil on 14 Apr 97 and have wound sets of pancakes for four more
of the 16 required coils.
During a site visit on 14 Apr 97, G. Biallas also discussed details
concerning manufacture of the 180-degree dipole coils and reverse-bend
dipoles with Magnet Enterprises International. They had generated
drawings for the tooling for potting the reverse-bend coils.
Process Equipment Company received the steel for these magnets'
cores and were programming the burning tables to create the blanks.
The coil contract for the injection-chicane dipoles was awarded
to Magnet Enterprises International, and winding details were
discussed with them. A local vendor, Master Machine, won the
contract for the cores.
Magnet measurements were done to determine the scaling of the
fringe field of the square Panofsky-style design for the trim
quadrupoles. Magnet measurements of a prototype Haimson coil
were also done. The Haimson coils have a 7% multipole content
while we need 1 %. Therefore they are out of the running for
our air-core correctors. We have identified two alternatives
thus far. We can make the coils with cosine-theta distributions
and flat ends for good field cut-off, and we have a line on some
that were made in the 1980's at Livermore Lab. At a minimum,
these are needed at the entrances and exits of the 180-degree
dipoles because they have to be able to run AC. We can also
try to use mu-metal (which saturates at 8 kG, while we need only
100 G fields for 42 MeV) to make window-frame magnet cores.
We have started to work on one of these in our Machine Shop.
The drawings of the sextupole coils were readied for signature
while their core drawings are in process.
The written procedure to weld the injection X-chamber is being
composed, and the fixturing is being prepared. Girders for the
back leg are nearly complete, while their QG-style quadrupoles
are being qualified on the Magnetic Test Stand.
Multipole measurements on the ten QB-style quadrupoles surplused
from CEBAF continue.
In the vacuum area, detailing of the chambers for the optical-chicane
dipoles is at 30%. The design of the reverse-bend chambers started
at Northrop-Grumman. The vacuum-pump order was placed. Design
of the chambers for the 180 degree dipoles neared completion.
The first rotating-coil-probe system for nulling the dipoles at
the switch points was received.
Regarding cryomodule fabrication:
Assembly of the first cryounit for the cryomodule continues unimpaired
and on schedule. Its helium vessel will be buttoned up and leak
checked today (18 Apr 97).
The second cavity pair was reprocessed this week, and it is now
in the Vertical Test Assembly awaiting retest.
Performance problems reported last week with regard to the third
cavity pair were traced to the power coupler. After correcting
the problem, we retested the pair successfully; it sustained >10
MV/m without field emission. This pair is presently awaiting
final leak check prior to accepting it for installation in its
Regarding electron-beam instrumentation and controls (I&C):
The requisition was signed for support stands for the beam-viewer
cameras. They each attach to the view ports and support both the
charge-coupled-device camera and the deflecting mirror.
The video multiplexers arrived, and the layout for the video distribution
is being finalized. At issue is which signals are to be viewed
The electronics filter banks and cable sets for the beam-position
monitors were ordered.
We set up a new "gateway" for I&C (and others) www
This gateway is found in section 5.11.2 of Jefferson Lab's FEL
Design work continues on the injector elements including the high
voltage power supply elbow, the gas handling system for the HVPS,
and the crane rail system in the injector area. Design work was
begun on the drive laser optical transport line.
Final drawings for the wiggler vacuum chamber and its support
were received. The machine shop has received the components
it needs to carry out hydrostatic pressing of the vacuum tubing
for the chamber. The final assembly drawings for the wiggler
girder are being checked. Almost all components needed for assembly
are in house. The wiggler has been assembled and shimmed for
the 12 mm gap at STI Optronics. The rms phase error is less
than 3 degrees, compared to the specification of 5 degrees.
This is the most critical parameter of the design. The inital
trajectory offset is on the order of 100 microns at 42 MeV and
the trajectory wander is less than 100 microns peak to peak.
This is excellent performance. The performance at 12 mm exceeds
all the specifications for the weak field gap. Measurements
of the weak field gap of around 15 mm are in progress at week's
end. The performance should be better than at the strong field
The custom 5-way crosses and flanges for the optical-cavity assemblies
have been faced off and cleaned for welding, which is planned
for next week. Development of the collimator continues to progress.
The optical-transport pipe in the FEL Facility's accelerator
enclosure has been welded, and it is awaiting delivery and installation
of the mirror cans. The Plexiglas used as laser shielding around
the mirror heat-sinking test stand is now installed.
A series of electron-beam simulations from photocathode to wiggler
was done to explore the implications of adjusting the charge per
bunch without changing the machine settings. The results were
generally favorable in that, for example, machine settings were
found that reflect a relative insensitivity of electron-beam properties
in going from first-light beam (60 pC per bunch) to full- power
beam (135 pC per bunch). A simulation of zero-current-beam propagation
with these machine settings pointed to some concerns about beam
degradation in this case. The concerns are pragmatic because
ghost pulses will be present, and they will behave as a zero-current
beam. As a consequence, some preliminary tracking studies through
the entire machine were done with this beam, and these tracking
studies suggest that no significant beam loss arises.
Another week of rapid progress on the Facility has come to pass.
Exterior panels were placed around the second floor, with the
exception of a small area near the injector where extra panels
are needed. Store-front work continues, with the cover on the
loading dock nearly complete. The framing in the lab areas was
completed early in the week, less some minor rework which will
be required. The framers then moved into the second-floor lobby
area and framed the bathrooms. Drywalling was completed in the
control room and optical-control room. The walls were sealed
and painted. The control-room window trim was cleaned and painted.
Drywall hanging began in several other key areas where its installation
holds up attachment of key components to the wall, including by
some electrical circuit panels, the ladder to the roof, etc.
A roof drain was moved to permit repositioning of the HVAC unit
by the control room. The water chiller for the HVAC was placed
and welded in to the piping in the equipment room. The electrical
transformer was turned on. Power was sent to both the upstairs
and downstairs switchgear, and the "real" lights were
turned on downstairs and partially upstairs. We now have main
power downstairs. Some modifications were made to the placement
of circuit breakers for the HVAC to get a large switchbox out
of the aisleway. Downstairs the welding and leak checking of
the cryogenic lines is proceeding well, and installation of stands
for the injector and injector dump has begun. Many additional
stands were marked out on the floor for bolt placement.
We decided this week to purchase the optical bench needed for
the Facility clean room. This keeps us from having to wait until
the ITS is shut down to move its optical bench. The cost of moving
was nearly one-third the cost of a new one, and the schedule
is helped substantially by the purchase.
We are pushing to get most of the cables pulled and penetrations
plugged by 1 Jul 97 and thereby eliminate the need for oxygen-deficiency-hazard
monitoring of the second floor. Dry nitrogen was added to the
services to be provided in the second-floor labs. We decided
to add a small platform outside the second-floor door to aid installation
of crane-delivered materials before the berm is complete. We
set a target date of 23 May 97 for interconnecting the Facility's
low- contamination-water systems with those of the nuclear-physics